# Neuro-Vision to Language: Image Reconstruction and Language enabled Interaction via Brain Recordings

### v2

@@ 4.1 Implementation Details
Dataset and Preprocessing: We utilized the Natural Scenes Dataset (NSD) [16], containing high-resolution 7Tesla fMRI scans and corresponding visual stimuli from COCO [29]. The dataset involved eight subjects, but analyses focused on the four (sub01, sub05, sub07) who completed all sessions. Modifications like cropping necessitated re-annotation of images using BLIP2 [30] for captions and DETR [31] for bounding boxes to maintain consistency. fMRI data was standardized to dimensions $83\times 104\times 81$ using trilinear interpolation and segmented into $14\times 14\times 14$ patches.
Architecture and Training: Our architecture integrates CLIP ViT-L/14 [14] and an Autoencoder KL [15] for image feature extraction, aligned with fMRI data processed through a 16-layer Transformer Encoder [32]. This setup employed two perceptrons ($f_{w_{c}}$, $f_{w_{v}}$) to align features with CLIP and VAE, respectively. Training involved a multi-stage approach where the LLM was initially frozen, followed by a fine-tuning stage for both the LLM and the Transformer Encoder. For visual reconstructions, the model utilized UnCLIP-2 [20] with $\beta$ set to 0.97, and concept localization was achieved using GradCAM [28]. For more details on the dataset and experimental setup, please refer to Appendices A,  B and  C. Additional experimental results can be found in Appendix D.
@@ 4.2 Captioning and Question Answering
Table 1: Quantitative analysis on brain captioning, detailed description, and complex reasoning tasks. The results are compared to other studies, with best, second, and third highlighted. The underline indicates the best result under the same conditions. denotes the results when using BLIP2-generated captions as the ground truth.
@@
Tab. 1 showcases the performance of our method on multimodal language tasks. With the introduction of LLMs, we have expanded the task forms to include not just brain captions, but also detailed descriptions and complex reasoning, as illustrated in Fig. 3. Our approach has demonstrated superior performance on the majority of metrics for the brain captioning task. Notably, our model can generalize across subjects without the necessity of training independent models for each subject.
Given the semantic mismatch between captions and images in the original NSD (Section 3.3). To accurately evaluate the brain captioning task, we reran the experiment using BLIP2 [30]-generated captions as ground truth. The results, shown in Tab. 1, show significant improvements when evaluated against BLIP2-generated captions, confirming the effectiveness of our model in the brain captioning task and the reasonableness of the task setting.
Beyond brain captioning, we have incorporated tasks for detailed description and complex reasoning. Our model also achieved the best performance on these two tasks, suggesting that it can generate not only simple captions but also detailed descriptions and perform complex reasoning. The model’s performance increases on complex reasoning tasks, possibly due to the richer semantic information in the questions, which our model captures more effectively. An ablation study was also conducted, revealing a noticeable performance drop in multimodal language tasks when the structural-preserving features of fMRI data extracted without ViT3d. This underlines the effectiveness of ViT3d and the capability of our model in multimodal tasks.
@@ 4.3 Visual Reconstruction
While our primary objective extends beyond mere visual decoding from fMRI data, visual reconstruction offers a tangible demonstration of a model’s comprehension of fMRI data. Therefore, we conducted visual reconstruction experiments and compared our results with those from other studies. The quantitative evaluation, highlights our method’s proficiency.
Table 2: Quantitative evaluation on visual reconstruction. Performance metrics are reported across different levels of features, with the best, second and third scores highlighted. The underline indicates the best result under the same conditions.
@@
Our Method $1$$284$370.370.370.370 $92.1\%$ $96.9\%$ $95.4\%$ $97.2\%$$612$282.282.282.282
w/o LLM $1$$283$369.369.369.369 $92.4\%$ $97.3\%$ $96.8\%$ $96.5\%$$647$273.273.273.273
w/o VAE feature $1$$093$263.263.263.263 $84.5\%$ $90.6\%$ $95.7\%$ $96.1\%$$691$391.391.391.391
w/o C_Subj $4$$267$356.356.356.356 $88.1\%$ $95.7\%$ $94.3\%$ $95.3\%$$627$327.327.327.327
w/o C_Subj & ViT3d $4$$194$273.273.273.273 $86.7\%$ $91.4\%$ $91.2\%$ $92.1\%$$713$403.403.403.403
@@
Our Method $1$$269$364.364.364.364 $90.7\%$ $94.6\%$ $94.6\%$ $95.8\%$$663$304.304.304.304
Tab. 2 showcases that our model competes with or surpasses traditional single-subject frameworks on several metrics. Particularly, it excels in high-level feature matching, demonstrating the model’s ability to effectively leverage Large Language Models for interpreting complex visual data. The robust performance across a wide array of visual stimuli confirms our model’s comprehensive understanding of fMRI data. Experiments without key components like LLM and VAE features highlight the significance of each element in our method, crucial for achieving state-of-the-art results. Moreover, we have conducted single-trial experiments, opting to use only the first visual stimulus, similar to the approach of MindEye, rather than averaging signals from three identical stimuli, which typically escalates data collection costs. Even under these more stringent conditions, our model shows only a slight decrease in performance, enhancing its feasibility for practical applications. Visual reconstruction examples are provided in Fig. 4, illustrating the effectiveness of our approach.
Figure 4: Visual reconstruction results showcasing the comparison between using the average signal from all trials (a) and using the first visual stimulus test (b).
@@
The balance between noise introduction and feature preservation in fMRI data image reconstruction is governed by the hyperparameter $\beta$. Fig. 5 presents a detailed ablation study on how different $\beta$ values impact various metrics. Fig. 5a shows that Pixel Correlation (PixCorr) peaks at intermediate $\beta$ values, indicating the optimal balance between injected noise and retained prior. The integration of the LLMs does not significantly influence low-level feature capture. In Fig. 5b, increasing $\beta$ enhances CLIP accuracy, with LLM integration having a substantial effect on capturing high-level features. Fig. 5c indicates that AlexNet layer 5 features, similar to CLIP features, effectively represent the similarity between reconstructed images and visual stimuli, capturing high-level features accurately. Additionally, Fig. 5d illustrates that both the Structural Similarity Index (SSIM) and CLIP scores benefit from optimally chosen $\beta$ values, with LLM integration enhancing overall image quality and semantic accuracy. Appropriately adjusting $\beta$ helps balance the representation of different feature levels in the reconstructed images. Fig. 6 provides examples of visual reconstructions at various $\beta$ values, demonstrating the model’s enhanced capabilities.
@@ 4.4 Concept Localization
To further our understanding of semantic concept localization within brain signals, we capitalized on the alignment between our fMRI encoder and CLIP features, developed during the training phase. Building on this, we devised a method to localize concepts within brain signals. Specifically, we first fine-tuned Language Models (LLMs) to extract the target concepts from natural language. These concepts, once encoded through the CLIP text encoder, served as targets for GradCAM, which facilitated the localization of the concept within brain signals. To enhance the precision of our localization, we trained three models with varying patch sizes (14, 12, 10) and utilized the penultimate layers of all models to extract semantic features. Fig. 7 illustrates the brain signal localization results for different semantic information, indicating our method’s capacity to discriminate the positions of various semantics within brain signals for the same visual stimulus.
@@
To validate the efficacy of our method, we conducted an ablation study on the semantic concepts. After localizing the concepts in the original brain signals, we zeroed out the signals in the identified region and performed feature extraction and visual reconstruction using the modified brain signals. As depicted in Fig. 8, the removal of neural activity in specific brain regions associated with certain semantic concepts resulted in the omission of the corresponding semantics in the visual reconstruction. This substantiates the validity of our concept localization method within brain signals and demonstrates our approach’s capacity for extracting and modifying semantic information in brain activity, which is pivotal for comprehending semantic information processing in the brain.
@@ 5 Conclusion
Our study has successfully developed and validated a novel brain decoding framework that leverages the capabilities of Vision Transformer 3D in conjunction with fMRI data, enhanced by the integration of LLMs. This approach has demonstrated a notable improvement in the reconstruction of visual stimuli from brain signals, offering a more precise and interpretable understanding of the underlying neural mechanisms. The experimental results confirmed the robustness of our model in performing various cognitive tasks, including captioning, question-answering, and visual reconstruction, all from single trial fMRI data. By enabling accurate localization of linguistic concepts within the brain, our work has potential applications in the development of brain-computer interfaces and advanced cognitive modeling.
Conclusively, this research contributes to the broader endeavor of decoding and interpreting brain activity, with significant implications for neuroscience and technology interface development. The fusion of advanced AI models with neuroimaging opens new avenues for exploring the intricacies of human cognition and the seamless integration of technology with neural processes.
@@ A.1 Natural Scenes Dataset
We conducted our experiments on the Natural Scenes Dataset (NSD) [16], which consists of high-resolution 7Tesla fMRI scans collected from eight healthy adult participants. Our analysis focused on the four subjects (subj01, subj02, subj05, and subj07) who completed all data collection sessions. Participants were exposed to thousands of natural scene images from the COCO dataset [29] during the sessions. However, the NSD required preprocessing to correct temporal resampling for slice timing differences and spatial interpolation to adjust for head motion and spatial distortion. We processed the scans in a 1.8-mm native volume space, particularly targeting the “nsdgeneral” region known for its high sensitivity to visual stimuli. This area, predominantly covering the posterior cortex, is crucial for targeted analysis of visual processing.
@@
A.2 Languate Extension
To ensure compatibility between fMRI data and Large Language Models (LLMs) and to enable instruction-following and diversified interactions, we extended the Natural Scenes Dataset (NSD) with natural language annotations. This extension includes four types of dialogues: brief descriptions, detailed descriptions, continuous dialogues, and complex reasoning tasks.
@@ A.2 Languate Extension
Furthermore, we constructed a multimodal fine-tuning dataset based on the dialogues generated. For each piece of fMRI data, we created corresponding language extensions. Commands for brief and detailed descriptions are illustrated in Tables 3 and  4, respectively. We randomly selected questions and corresponding answers from these to construct Q&A pairs, enhancing the model’s ability to engage in meaningful dialogue based on the visual content.
@@ Appendix C Architecture and Training Details
CLIP ViT-L/14 [14] and Autoencoder KL [15] were used as feature extractors for images, aligning with fMRI data. For the fMRI data, we employed a 16-layer Transformer Encoder [32] with a hidden size of 768 to extract features, using the class token from the last layer as the output. Two two-layer perceptrons with a hidden dimension of 1024, $f_{w_{c}}$ and $f_{w_{v}}$, were used to align with CLIP and VAE features, respectively, with the hyperparameter $\alpha$ set to $1/64$. The learning rate was set to $5\times 10^{-4}$, training for $30$ epochs. For multimodal interaction, the aforementioned Transformer Encoder was frozen, using its second-to-last layer’s hidden states as the fMRI token, processed through a two-layer perceptron $f_{w_{t}}$, to interact with Vicuna-13B [39]. Training was divided into two stages: initially, the LLM was frozen, tuning only $f_{w_{t}}$, and in the second stage, both the LLM and $f_{w_{t}}$ were fine-tuned simultaneously. The learning rate was set to $2\times 10^{-5}$, training for one epoch. In the visual reconstruction phase, the LLM called upon UnCLIP-2 [20] for image reconstruction, with the hyperparameter $\beta$ set to 0.97. For concept location, the LLM first extracted keywords and then localized them using GradCAM [28], visualizing the results.

### v4

@@ 4.1 Implementation Details
Dataset and Preprocessing: We utilized the Natural Scenes Dataset (NSD) [16], containing high-resolution 7Tesla fMRI scans and corresponding visual stimuli from COCO [34]. The dataset involved eight subjects, but analyses focused on the four (subj01, subj02, subj05 and subj07) who completed all sessions. Modifications like cropping necessitated re-annotation of images using BLIP2 [36] for captions and DETR [37] for bounding boxes to maintain consistency. fMRI data was standardized to dimensions $83\times 104\times 81$ using trilinear interpolation and segmented into $14\times 14\times 14$ patches.
Architecture and Training: Our architecture integrates CLIP ViT-L/14 [14] and an AutoencoderKL [15] for image feature extraction, aligned with fMRI data processed through a $16$-layer Transformer Encoder [28]. This setup employed two perceptrons ($f_{w_{c}}$, $f_{w_{v}}$) to align features with CLIP and VAE, respectively. Training involved a multi-stage approach where the LLM was initially frozen, followed by a fine-tuning stage for both the LLM and the Transformer Encoder. For visual reconstructions, the model utilized UnCLIP-2 [20] with $\beta$ set to $0.93$, and concept localization was achieved using GradCAM [35]. For more details on the dataset and experimental setup, please refer to Appendices A,  B, and  C. Additional experimental results can be found in Appendix D.
@@ 4.2 Captioning and Question Answering
Table 1: Quantitative analysis of brain captioning, detailed descriptions, and complex reasoning tasks. Some results are derived from UMBRAE [24]. Results are compared to those from other studies, with best, second, and third highlighted. Underline indicates the best result under identical conditions, while denotes results obtained using BLIP2-generated captions as ground truth.
@@
Tab. 1 shows the performance of our method on multimodal language tasks. With the introduction of LLMs, we have expanded the task forms to include brain captions, detailed descriptions, and complex reasoning, as illustrated in Fig. 3. Our approach has demonstrated superior performance on the majority of metrics for the brain captioning task. Notably, our model can generalize across subjects without the need to train separate models for each subject or introduce subject-specific parameters.
Given the semantic mismatch between captions and images in the original NSD (Section 3.3), we reran the experiment using BLIP2 [36]-generated captions as ground truth. The results, shown in Tab. 1, show significant improvements when evaluated against BLIP2-generated captions, confirming the effectiveness of our model in the brain captioning task and the reasonableness of the task setting.
Beyond brain captioning, we have incorporated tasks for detailed description and complex reasoning. Our model also achieved the best performance on these two tasks, suggesting that it can generate not only simple captions but also detailed descriptions and perform complex reasoning. The model’s performance increases on complex reasoning tasks, possibly due to the richer semantic information in the questions, which our model captures more effectively. An ablation study was also conducted, revealing a noticeable performance drop in multimodal language tasks when the structural-preserving features of fMRI data were not extracted using ViT3D. Instead, the fMRI data were flattened and patched, similar to methods used in other literature, while maintaining the same fMRI feature extractor structure. This underlines the effectiveness of ViT3D and the capability of our model in multimodal tasks.
@@ 4.3 Visual Reconstruction
While our primary objective extends beyond mere visual decoding from fMRI data, visual reconstruction offers a tangible demonstration of a model’s comprehension of fMRI data. Therefore, we conducted visual reconstruction experiments and compared our results with those from other studies. The quantitative evaluation highlights our method’s proficiency.
Table 2: Quantitative evaluation on visual reconstruction. Performance metrics are reported across different levels of features, with the best, second and third scores highlighted. The underline indicates the best result under the same conditions. Our method achieves state-of-the-art results using a single model trained across subjects (# Models = 1).
@@
MindBridge [25] $1$$151$263.263.263.263 $87.7\%$ $95.5\%$ $92.4\%$ $94.7\%$$712$418.418.418.418
UMBRAE [24] $1$$283$328.328.328.328 $93.9\%$ $96.7\%$ $93.4\%$ $94.1\%$$700$393.393.393.393
Our Method $1$$265$357.357.357.357 $93.1\%$ $97.1\%$ $96.8\%$ $97.5\%$$633$321.321.321.321
w/o LLM $1$$263$369.369.369.369 $92.0\%$ $97.1\%$ $94.2\%$ $96.1\%$$680$328.328.328.328
w/o VAE feature $1$$093$263.263.263.263 $84.5\%$ $90.6\%$ $93.6\%$ $95.7\%$$684$398.398.398.398
w/o C_Subj $4$$241$356.356.356.356 $88.1\%$ $95.7\%$ $92.1\%$ $95.1\%$$631$347.347.347.347
w/o C_Subj & ViT3D $4$$164$273.273.273.273 $86.7\%$ $91.4\%$ $89.3\%$ $91.8\%$$731$417.417.417.417
@@
Our Method $1$$257$336.336.336.336 $91.2\%$ $96.3\%$ $94.6\%$ $95.3\%$$671$324.324.324.324
Tab. 2 showcases that our model competes with or surpasses traditional subject-specific frameworks on several metrics. Notably, it excels in high-level feature matching, demonstrating the model’s ability to effectively leverage LLMs for interpreting complex visual data. The robust performance across various visual stimuli confirms our model’s comprehensive understanding of fMRI data. Experiments without key components like LLM and VAE features highlight the significance of each element in our method, which is crucial for achieving state-of-the-art results. Moreover, we have conducted single-trial experiments, opting to use only the first visual stimulus, similar to the approach of MindEye [6], rather than averaging signals from three identical stimuli, which typically escalates data collection costs. Even under these more stringent conditions, our model shows only a slight decrease in performance, enhancing its feasibility for practical applications. Visual reconstruction examples are provided in Fig. 4, illustrating the effectiveness of our approach.
Figure 4: Visual reconstruction results showcasing the comparison between (a) using the average signal from all trials and (b) using the first visual stimulus.
@@
The balance between noise introduction and feature preservation in fMRI data visual reconstruction is governed by the hyperparameter $\beta$. Fig. 5 presents a detailed ablation study on how different $\beta$ values impact various metrics. Fig. 5(a) shows that Pixel Correlation (PixCorr) peaks at intermediate $\beta$ values, indicating the optimal balance between injected noise and retained prior. The integration of the LLMs does not significantly influence low-level feature capture. In Fig. 5(b), increasing $\beta$ enhances CLIP accuracy, with LLM integration having a substantial effect on capturing high-level features. Fig. 5(c) indicates the features of the $5$ th layer of AlexNet, similar to CLIP features, effectively represent the similarity between reconstructed images and visual stimuli, capturing high-level features accurately. Additionally, Fig. 5(d) illustrates that both the Structural Similarity Index (SSIM) and CLIP scores benefit from optimally chosen $\beta$ values, with LLM integration enhancing overall image quality and semantic accuracy. Appropriately adjusting $\beta$ helps balance the representation of different feature levels in the reconstructed images. Fig. 6 provides examples of visual reconstructions at various $\beta$ values, demonstrating the model’s enhanced capabilities.
@@ 4.4 Concept Localization
To further our understanding of semantic concept localization within brain signals, we capitalized on the alignment between our fMRI encoder and CLIP features, which were developed during the training phase. Building on this, we devised a method to localize concepts within brain signals. Specifically, we first fine-tuned Language Models (LLMs) to extract the target concepts from natural language. These concepts, once encoded through the CLIP text encoder, served as targets for GradCAM, which facilitated the localization of the concept within brain signals. To enhance the precision of our localization, we trained three models with varying patch sizes $(14,12,10)$ and utilized the penultimate layers of all models to extract semantic features. Fig. 7 illustrates the brain signal localization results for different semantic information, indicating our method’s capacity to discriminate the positions of various semantics within brain signals for the same visual stimulus.
@@
To validate the efficacy of our method, we conducted an ablation study on the semantic concepts. After localizing the concepts in the original brain signals, we zeroed out the signals in the identified voxels and performed feature extraction and visual reconstruction using the modified brain signals. As depicted in Fig. 8, the removal of neural activity in specific brain regions associated with certain semantic concepts resulted in the omission of the corresponding semantics in the visual reconstruction. This substantiates the validity of our concept localization method within brain signals and demonstrates our approach’s capacity for extracting and modifying semantic information in brain activity, which is pivotal for comprehending semantic information processing in the brain.
@@ 5 Conclusion
Our study has successfully developed and validated a novel brain decoding framework that leverages the capabilities of Vision Transformer 3D in conjunction with fMRI data, enhanced by the integration of LLMs. This approach has demonstrated a notable improvement in the reconstruction of visual stimuli from brain signals, offering a more precise and interpretable understanding of the underlying neural mechanisms. The experimental results confirmed the robustness of our model in performing various cognitive tasks, including captioning, question-answering, and visual reconstruction, all from single-trial fMRI data. By enabling accurate localization of linguistic concepts within the brain, our work has potential applications in developing brain-computer interfaces and advanced cognitive modeling. Conclusively, this research contributes to the broader endeavor of decoding and interpreting brain activity, with significant implications for neuroscience and technology interface development. The fusion of advanced AI models with neuroimaging opens new avenues for exploring the intricacies of human cognition and the seamless integration of technology with neural processes.
Acknowledgments and Disclosure of Funding
This research was financially supported by a funding from Institute of Automation, Chinese Academy of Sciences (Grant No. E411230101).
@@ A.1 Natural Scenes Dataset
We conducted our experiments on the Natural Scenes Dataset (NSD) [16], which consists of high-resolution 7Tesla fMRI scans collected from eight healthy adult participants. Our analysis focused on the four subjects (subj01, subj02, subj05, and subj07) who completed all data collection sessions. Participants were exposed to thousands of natural scene images from the COCO dataset [34] during the sessions. However, the NSD required preprocessing to correct temporal resampling for slice timing differences and spatial interpolation to adjust for head motion and spatial distortion. We processed the scans in a 1.8-mm native volume space, particularly targeting the "nsdgeneral" region known for its high sensitivity to visual stimuli. This area, predominantly covering the posterior cortex, is crucial for targeted analysis of visual processing.
@@
A.2 Language Extension
To ensure compatibility between fMRI data and Large Language Models (LLMs) and to enable instruction-following and diversified interactions, we extended the Natural Scenes Dataset (NSD) with natural language annotations. This extension includes seven types of dialogues: brief descriptions, detailed descriptions, continuous dialogues, complex reasoning tasks, instruction reconstruction, and concept localization.
@@ A.2 Language Extension
Furthermore, we constructed a multimodal fine-tuning dataset based on the generated dialogues. For each piece of fMRI data, we created corresponding language extensions. Commands for brief and detailed descriptions are illustrated in Tab. 3 and  4, respectively. We randomly selected questions and corresponding answers from these to construct Q&A pairs, enhancing the model’s ability to engage in meaningful dialogue based on the visual content. For continuous dialogues and complex reasoning, we used dialogues generated by GPT-3.5-turbo-0125. For instruction reconstruction, the commands are shown in Tab. 5, which aim to have the model generate detailed descriptions of visual stimuli using concise expressions. For concept localization, the commands are shown in Tab. 6, used to extract concepts mentioned in the prompts and visualize the model’s attention using grad-CAM [35].
@@
"Provide the corresponding Stable Diffusion prompts for the image.",
Table 5: The list of instructions for instruction reconstruction.
"Locating the concept of "<object>",
Table 6: The list of instructions for concept localization.
@@ Appendix C Architecture and Training Details
CLIP ViT-L/14 [14] and Autoencoder KL [15] were used as feature extractors for images, aligning with fMRI data. For the fMRI data, we employed a 16-layer Transformer Encoder [28] with a hidden size of 768 to extract features, using the class token from the last layer as the output. Two two-layer perceptrons with a hidden dimension of 1024, $f_{w_{c}}$ and $f_{w_{v}}$, were used to align with CLIP and VAE features, respectively, with the hyperparameter $\alpha$ set to $1/64$. The learning rate was set to $5\times 10^{-4}$, training for $30$ epochs. For multimodal interaction, the aforementioned Transformer Encoder was frozen, using its second-to-last layer’s hidden states as the fMRI token, processed through a two-layer perceptron $f_{w_{t}}$, to interact with Llama-3-8B [44]. Training was divided into two stages: initially, the LLM was frozen, tuning only $f_{w_{t}}$, and in the second stage, both the LLM and $f_{w_{t}}$ were fine-tuned simultaneously. The learning rate was set to $2\times 10^{-5}$, training for one epoch. In the visual reconstruction phase, the LLM called upon UnCLIP-2 [20] for visual reconstruction, with the hyperparameter $\beta$ set to $0.93$. For concept location, the LLM first extracted keywords and then localized them using GradCAM [35], visualizing the results.
@@ D.1 Visual Reconstruction
Table 7: Performance comparison on visual reconstruction tasks when using different LLMs and different instructions. The underline indicates the best result under the same conditions.
Method Instruction Low-Level High-Level
w/o LLM - .$263$369.369.369.369 $92.0\%$ $97.1\%$ $94.2\%$ $96.1\%$$680$328.328.328.328
w/ Vicuna-13B briefly descriptions .$259$351.351.351.351 $91.5\%$ $96.6\%$ $95.1\%$ $96.7\%$$641$337.337.337.337
w/ Vicuna-13B instruction reconstruction .$257$361.361.361.361 $92.1\%$ $96.9\%$ $96.2\%$ $97.1\%$$628$331.331.331.331
w/ Llama-3-8B briefly descriptions .$261$354.354.354.354 $92.7\%$ $96.8\%$ $96.4\%$ $97.0\%$$637$327.327.327.327
w/ Llama-3-8B instruction reconstruction .$265$357.357.357.357 $93.1\%$ $97.1\%$ $96.8\%$ $97.5\%$$633$321.321.321.321
Tab. 7 shows the performance comparison of visual reconstruction tasks when using different large language models (LLMs) and different generation instructions. We compared the two models Vicuna-13B [50] and Llama-3-8B [44]. The results indicate that our method is compatible with various LLMs and performs better in visual reconstruction tasks when using more powerful LLMs. When using Llama-3-8B, the best performance was achieved across the majority of metrics. Additionally, we explored the impact of different generation instructions on the visual reconstruction task. The results show that using instructions specifically designed for visual reconstruction can improve the performance of the visual reconstruction task. We believe this is because these instructions enable the LLMs to generate more detailed descriptions of visual stimuli, thereby enhancing the performance of the visual reconstruction task.